A clock's minute hand is 10cm long. What area it will cover from 9:00am to 9:35am?
Answer by Cromlix(4381)

360 degrees on a clock face.
Each 5 mins = 30 degrees.
So, 9.00 - 9.35 = 7*30 = 210 degrees.
210/360 = area/Pi*r^2
210/360 = area/Pi*10^2
Area = (210*Pi*100)/360
Area = 183.3cm^2
Hope this helps.
:-)